# Multi-scale Protein Language Model for Unified Molecular Modeling

### v2

@@ Abstract
Protein language models have demonstrated significant potential in the field of protein engineering. However, current protein language models primarily operate at the residue scale, which limits their ability to provide information at the atom level. This limitation prevents us from fully exploiting the capabilities of protein language models for applications involving both proteins and small molecules. In this paper, we propose ESM-AA (ESM All-Atom), a novel approach that enables atom-scale and residue-scale unified molecular modeling. ESM-AA achieves this by pre-training on multi-scale code-switch protein sequences and utilizing a multi-scale position encoding to capture relationships among residues and atoms. Experimental results indicate that ESM-AA surpasses previous methods in protein-molecule tasks, demonstrating the full utilization of protein language models. Further investigations reveal that through unified molecular modeling, ESM-AA not only gains molecular knowledge but also retains its understanding of proteins.
@@ 2.2 Multi-scale Pre-training
In this section, we elaborate how to create a code-switch protein sequence $\bar{X}$ and implement the pre-training tasks, i.e., MLM and PDR, on it (Figure 2).
@@ 2.4 Integrating Multi-scale PE into Transformer
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.623 0.548 0.742 90.6% 0.750 0.943
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.599 0.566 0.753 91.8% 0.781 0.954
@@
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.248 0.889 0.668
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.191 0.906 0.759
@@ Results
Table 3: Performance comparison on the Contact Prediction (CP) task and Secondary Structure Prediction (SSP) task. ESM-AA has similar performance to ESM-2 indicating it preserves the local and global understanding of proteins originally presented in ESM-2.
Method CP (P@L/5$\uparrow$ SSP (ACC)$\uparrow$
Short Range Medium Range Long Range SS3-casp12 SS8-casp12
TAPE 38M 0.46 0.33 0.25 0.71 0.59
ResNet 38M 0.46 0.35 0.17 0.72 0.58
ESM-2 35M 0.46 0.45 0.49 0.74 0.61
ESM-AA 35M 0.48 0.45 0.48 0.74 0.60
@@ 3.4 ESM-AA Preserves the Strong Ability of Protein Understanding
Because ESM-AA is developed based on existing PLMs, we would like to determine whether it still preserves a thorough understanding of proteins. Specifically, we follow TAPE (Rao et al., 2019) and use the tasks secondary structure prediction and contact prediction to test the ability of protein pre-training models in protein structure understanding. For secondary structure prediction, models must grasp the local protein structure, such as helices and strands. For contact prediction, models need a comprehensive understanding of proteins at a global level. Notably, both ESM-AA and baseline methods have exactly the same input (pure residue sequence) for these two tasks. For more details of the fine-tuning and baselines (datasets, framework, and hyperparameters), readers can find them in Appendix B.2.
We report the results of contact prediction and secondary structure prediction in Table 3. While ESM-AA may not achieve the best performance among the compared methods, the tables demonstrate that it performs similarly to ESM-2 in both secondary structure prediction and contact prediction. This indicates that ESM-AA does not sacrifice its understanding of proteins. Promisingly, ESM-AA can achieve improved protein understanding by initializing its parameters with a larger ESM-2. More results on these two protein tasks are shown in Appendix C, specifically in Table 7 and Table 8.
@@ Appendix B Fine-tuning Details
Table 6: The statistics of downstream datasets in one table. ESAR: Enzyme-Substrate Affinity Regression, DTAR: Drug-Target Affinity Regression, ESPC: Enzyme-Substrate Pair Classification, SSP: Secondary Structure Prediction, CP: Contact Prediction, MPR: Molecular Property Regression, MPC: Molecular Property Classification.
@@
Task ESAR DTAR ESPC SSP CP MPR MPC
@@ B.2 Fine-tuning Details of Protein Tasks
Following TAPE’s protocol (Rao et al., 2019), we evaluate ESM-AA on secondary structure prediction and contact prediction tasks. Specifically, for secondary structure prediction, we use data from Klausen et al. (2019) as training and validation sets and use CB513 (Cuff & Barton, 1999), CASP12 (Moult et al., 2018), and TS115 (Yang et al., 2018) as test sets. The training and validation sets are filtered at the 25% sequence identity threshold with these test tests. The final training, validation and three test sets have 8678, 2170, 513, 21, 115 protein sequences, respectively. For contact prediction tasks, we use training, validation, and test sets from ProteinNet (AlQuraishi, 2019) with training and validation sets filtered at the 30% sequence identity threshold. For a fair comparison, we also remove the test data that appears in the pre-training data, and the proportion of this part of the data is less than 4‰. The final training, validation, and test sets have 20, 24, 13945 protein sequences.
@@
As suggested by TAPE, for both protein-only tasks, we use ESM-AA as the protein encoder. When doing secondary structure prediction, we use a linear output layer to predict the secondary structure which each residue belongs to. When handling the contact prediction task, we use the attention from the last layer as features and then use a linear layer to predict whether these two residues have contact or not. Notably, both input of these two tasks is only protein sequences without structural information. Therefore, when using ESM-AA to handle these two tasks, we turn off the unzip.
@@ B.3 Fine-tuning Details of Molecule Tasks
Appendix C More Experiment Results
Table 7 and Table 8 display the completed experimental results for the contact prediction and secondary structure prediction tasks. Table 9 shows the experiment results of both molecular property classification and regression tasks.
Table 7: Performance comparison on the contact prediction task.
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
TAPE 38M 0.28 0.35 0.46 0.19 0.25 0.33 0.17 0.20 0.25
ResNet 38M 0.25 0.34 0.46 0.18 0.25 0.35 0.10 0.13 0.17
ESM-2 35M 0.20 0.29 0.46 0.22 0.32 0.45 0.30 0.39 0.49
ESM-AA 35M 0.21 0.31 0.48 0.23 0.32 0.45 0.29 0.38 0.48
Method SS3(ACC) $\uparrow$ SS8(ACC) $\uparrow$
cb513 ts115 casp12 cb513 ts115 casp12
TAPE 38M 0.73 0.77 0.71 0.59 0.64 0.59
ResNet 38M 0.75 0.78 0.72 0.58 0.64 0.58
ESM-2 35M 0.80 0.82 0.74 0.65 0.70 0.61
ESM-AA 35M 0.79 0.81 0.74 0.63 0.69 0.60

### v3

# ESM All-Atom: Multi-scale Protein Language Model for Unified Molecular Modeling

@@ Abstract
Protein language models have demonstrated significant potential in the field of protein engineering. However, current protein language models primarily operate at the residue scale, which limits their ability to provide information at the atom level. This limitation prevents us from fully exploiting the capabilities of protein language models for applications involving both proteins and small molecules. In this paper, we propose ESM-AA (ESM All-Atom), a novel approach that enables atom-scale and residue-scale unified molecular modeling. ESM-AA achieves this by pre-training on multi-scale code-switch protein sequences and utilizing a multi-scale position encoding to capture relationships among residues and atoms. Experimental results indicate that ESM-AA surpasses previous methods in protein-molecule tasks, demonstrating the full utilization of protein language models. Further investigations reveal that through unified molecular modeling, ESM-AA not only gains molecular knowledge but also retains its understanding of proteins.  [^1] The source codes of ESM-AA  are publicly released at https://github.com/zhengkangjie/ESM-AA.
@@ 2.2 Multi-scale Pre-training
In this section, we elaborate how to create a code-switch protein sequence $\bar{X}$ and implement the pre-training tasks, i.e., Masked Language Modeling (MLM) and Pair-wise Distance Recovery (PDR), on it (Figure 2).
@@ 2.4 Integrating Multi-scale PE into Transformer
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.620 0.551 0.744 90.4% 0.743 0.949
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.607 0.560 0.752 92.3% 0.797 0.957
@@
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.243 0.890 0.678
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.196 0.903 0.752
@@ 3.4 ESM-AA Preserves the Strong Ability of Protein Understanding
Method SS3(ACC) $\uparrow$ SS8(ACC) $\uparrow$
cb513 ts115 casp12 cb513 ts115 casp12
TAPE 38M 0.73 0.77 0.71 0.59 0.64 0.59
ResNet 38M 0.75 0.78 0.72 0.58 0.64 0.58
ESM-2 35M 0.80 0.82 0.74 0.65 0.70 0.61
ESM-AA 35M 0.79 0.81 0.74 0.63 0.69 0.60
Table 5: Performance comparison on the unsupervised contact prediction task.
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
TAPE 92M 0.10 0.12 0.16 0.10 0.13 0.17 0.11 0.14 0.18
ESM-1 43M 0.11 0.13 0.16 0.12 0.15 0.19 0.13 0.17 0.22
ESM-2 35M 0.20 0.29 0.46 0.22 0.32 0.45 0.30 0.39 0.49
ESM-AA 35M 0.21 0.31 0.48 0.23 0.32 0.45 0.29 0.38 0.48
Because ESM-AA is developed based on existing PLMs, we would like to determine whether it still preserves a thorough understanding of proteins. Specifically, we follow TAPE (Rao et al., 2019), ESM(Rao et al., 2020) and use the tasks secondary structure prediction and unsupervised contact prediction to test the ability of protein pre-training models in protein structure understanding. For secondary structure prediction, models must grasp the local protein structure, such as helices and strands. For unsupervised contact prediction, models need a comprehensive understanding of proteins at a global level. Notably, both ESM-AA and baseline methods have exactly the same input (pure residue sequence) for these two tasks. For more details of the fine-tuning and baselines (datasets, framework, and hyperparameters), readers can find them in Appendix B.2.
We report the results of secondary structure prediction and unsupervised contact prediction in Table 4 and Table 5. While ESM-AA may not achieve the best performance among the compared methods, the tables demonstrate that it performs similarly to ESM-2 in both secondary structure prediction and contact prediction. This indicates that ESM-AA does not sacrifice its understanding of proteins. Promisingly, ESM-AA can achieve improved protein understanding by initializing its parameters with a larger ESM-2.
@@ Appendix B Fine-tuning Details
Table 7: The statistics of downstream datasets in one table. ESAR: Enzyme-Substrate Affinity Regression, DTAR: Drug-Target Affinity Regression, ESPC: Enzyme-Substrate Pair Classification, SSP: Secondary Structure Prediction, UCP: Unsupervised Contact Prediction, MPR: Molecular Property Regression, MPC: Molecular Property Classification.
@@
Task ESAR DTAR ESPC SSP UCP MPR MPC
@@ B.2 Fine-tuning Details of Protein Tasks
Following TAPE’s protocol (Rao et al., 2019), we evaluate ESM-AA on secondary structure prediction. Specifically, for secondary structure prediction, we use data from Klausen et al. (2019) as training and validation sets and use CB513 (Cuff & Barton, 1999), CASP12 (Moult et al., 2018), and TS115 (Yang et al., 2018) as test sets. The training and validation sets are filtered at the 25% sequence identity threshold with these test tests. The final training, validation and three test sets have 8678, 2170, 513, 21, 115 protein sequences, respectively. Following ESM’s protocol (Rao et al., 2020), we use training, validation, and test sets from ProteinNet (AlQuraishi, 2019) with training and validation sets filtered at the 30% sequence identity threshold for unsupervised contact prediction tasks. For a fair comparison, we also remove the test data that appears in the pre-training data, and the proportion of this part of the data is less than 4‰. The final training, validation, and test sets have 20, 24, 13945 protein sequences.
@@
As suggested by TAPE, for both protein-only tasks, we use ESM-AA as the protein encoder. When doing secondary structure prediction, we use a linear output layer to predict the secondary structure which each residue belongs to. When handling the unsupervised contact prediction task, we use the attention from each layer and head are independently symmetrized and corrected with APC (Dunn et al., 2008) as features and then use a linear layer to predict whether these two residues have contact or not. Notably, both input of these two tasks is only protein sequences without structural information. Therefore, when using ESM-AA to handle these two tasks, we turn off the unzip.
@@ B.3 Fine-tuning Details of Molecule Tasks
Appendix C More Experiment Results on Molecular Tasks
Table 8 shows the experiment results of both molecular property classification and regression tasks.
@@ Appendix G More Ablation Results
Ablation on Protein-only Tasks
We have tested the performance of model ablation experiments on the Contact Prediction task. And the results are shown in Table 12. The absence of MLM Loss will have the most significant adverse effect on the model’s performance. This is because MLM at the amino acid scale is the primary means for the model to learn semantic information about proteins. PDR (pair-wise distance recovery) performed within individual residues does not assist the model in learning global semantic information about proteins. Removing the MLM Loss will result in the model being unable to learn meaningful protein representations from the data. Removing Residue Scale Position Encoding (w/o RSPE) as well as removing protein data (w/o Protein Data) will also significantly impact the model’s ability to learn protein representations. This demonstrates the necessity of Residue Scale Position Encoding. The presence or absence of the Unzip Operation does not significantly affect the model’s performance on tasks such as Contact Prediction, where sequences are used as input. This indicates that the protein’s local structural information introduced by the Unzip Operation does not directly impact the model’s performance.
Table 12: The scaling experimental results on contact map prediction.
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
w/o ASPE 0.19 0.28 0.43 0.20 0.28 0.40 0.25 0.33 0.42
w/o RSPE 0.05 0.06 0.05 0.04 0.04 0.04 0.02 0.02 0.03
w/o MLM Loss 0.04 0.03 0.03 0.03 0.03 0.02 0.03 0.03 0.03
w/o PDR Loss 0.11 0.14 0.18 0.10 0.12 0.15 0.10 0.12 0.16
w/o Protein Data 0.05 0.05 0.04 0.04 0.03 0.03 0.04 0.04 0.03
w/o Molecule Data 0.20 0.29 0.46 0.22 0.31 0.45 0.29 0.38 0.48
w/o Unzip Operation 0.20 0.30 0.46 0.22 0.32 0.45 0.30 0.38 0.49
ESM-AA 0.20 0.30 0.46 0.22 0.32 0.45 0.29 0.38 0.48
Ablation on Molecule-only Tasks
Ablation studies on molecule-only tasks prove the importance of the Unzip operation in learning good molecular representations. These results are shown in Table 13. From the results, it can be seen that the absence of ASPE and Unzip Operation will have the most significant adverse effect on the model’s performance. This is because ASPE serves as the unique identifier for the model to distinguish between different atoms, while the Unzip Operation can introduce diverse residue structural information to the model. Both of these are key in improving the model’s modeling of atomic-scale information. Even though the model’s performance declines after removing molecular training data, it still maintains a relatively high level. This is because the Unzip operation unfolds some residues into atomic-scale information, allowing the model to learn important atomic-scale semantic representations.
Table 13: Ablation studies on molecule-only tasks.
Method BACE $\uparrow$ BBBP $\uparrow$ MUV $\uparrow$ HIV $\uparrow$
w/o ASPE 73.99 66.41 48.93 75.22
w/o RSPE 75.15 70.31 62.62 75.81
w/o MLM Loss 79.15 65.09 65.91 76.46
w/o PDR Loss 76.42 64.23 61.22 74.55
w/o Protein Data 74.9 1 66.9 71.49 78.96
w/o Molecule Data 77.8 69.12 72.31 72.05
w/o Unzip Operation 61.57 65.34 59.59 77.21
ESM-AA 83.5 70.1 76.2 77.3
Appendix H Scaling law of ESM-AA
We also conducted scaling experiments at various scales and concluded that ESM-AAadheres to the scaling law. However, the upper limit of its capabilities on single-modal data restricts the scale-up of ESM-AAat this stage. Here are the detailed results.
H.1 Scaling ESM-AA  from 8M to 35M.
Scaling ESM-AA  from 8M to 35M significantly improves performance on protein-molecule tasks. We test the model’s performance on protein-molecule tasks(results are shown in Table 14), protein-only tasks(results are shown in Table 16), and molecule-only tasks(results are shown in Table 15), and the experimental results demonstrate that the 35M-scale ESM-AA  significantly outperforms the 8M-scale ESM-AA .
H.2 Scaling ESM-AA  from 35M to 150M.
Scaling ESM-AA  from 35M to 150M significantly improves performance on protein-only tasks but does not result in significant performance gains for protein-molecule tasks. To further analyze the reasons behind the above phenomenon, we conducted tests on the scaled-up model’s performance on molecule-only tasks.
On molecule-only tasks, the performance of the 150M model showed minimal improvement. To confirm the lack of improvement in molecular performance due to scaling up, we also trained a 150M-sized Uni-Mol model. Results in Table 15 showed that the 150M Uni-Mol model exhibited almost no performance growth compared to the 47M Uni-Mol model. Further investigation into current 3D molecular representation learning models revealed that mainstream models are generally smaller than 50M, and there has been no work observed to date that extends 3D molecular representation learning models beyond 100M in size. This indicates that scaling up does not significantly benefit molecular performance for current models and datasets. On protein-only tasks, ESM-AA  150M can significantly outperform ESM-AA  35M. In summary, the bottleneck of scaling up lies in the model’s ability to learn molecular representations. In the future, we will further explore how to successfully scale up molecular representation learning models. We will also incorporate these discussions and results into the paper later on.
Table 14: The scaling results on enzyme-substrate affinity regression task.
Protein Molecule ESAR
ESM-2 35M Uni-Mol 48M 0.642 0.536 0.733
ESM-AA 8M ESM-AA 8M 0.618 0.552 0.734
ESM-AA 35M ESM-AA 35M 0.599 0.566 0.753
ESM-AA 150M ESM-AA 150M 0.626 0.547 0.741
Table 15: The scaling experimental results on molecule-only tasks.
Method BACE $\uparrow$ BBBP $\uparrow$ MUV $\uparrow$ HIV $\uparrow$
ESM-AA 8M 75.9 67.3 70.7 74.3
ESM-AA 35M 83.5 70.2 76.2 77.3
ESM-AA 150M 83.8 71.2 71.8 78.8
Uni-Mol 48M 83.2 71.5 72.0 78.3
Uni-Mol 150M 83.9 71.4 71.1 78.5
Table 16: The scaling experimental results on contact map prediction.
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
ESM-AA  8M 0.15 0.21 0.31 0.14 0.19 0.26 0.15 0.20 0.27
ESM-AA  35M 0.21 0.31 0.48 0.23 0.32 0.45 0.29 0.38 0.48
ESM-AA  150M 0.24 0.38 0.60 0.29 0.42 0.61 0.43 0.53 0.69
ESM-2 35M 0.20 0.29 0.46 0.22 0.32 0.45 0.30 0.39 0.49
ESM-2 150M 0.24 0.38 0.59 0.28 0.42 0.61 0.43 0.55 0.68
Appendix I How to Choose the Unzip Proportion
The reason for choosing an unzip proportion of 1% is a balanced decision considering both performance and training cost. We have validated that the 1% unzip ratio parameter is a relatively good choice through some experiments.
A high unzip ratio will incur high training costs. We find that as the unzip proportion increases, the number of tokens in the data significantly increases, as does the length of protein sequences. This leads to an increase in training cost. Therefore, choosing a proportion that is too large is not conducive to completing the training process with limited computational resourcs. We have experimentally verified and determined the optimal unzip ratio selection. We compared the model performance under three scenarios: unzip ratio of 0, 1%, and 5%. We found that when the unzip ratio is set to 1%, the model exhibits the best performance in the protein-molecule task(shown in Table 17). When the unzip proportion is too small, the model’s performance also decreases, especially in terms of molecular representation learning performance (shown in Table 17). This is because when more residues are unfolded, the model can obtain more atomic-scale training data, which is more conducive to learning unified semantic representations. Taking these factors into consideration, we ultimately choose 0.01 as the unzip proportion. At this proportion, approximately 8% of the tokens in the final protein sequence are atomic-scale tokens. This falls well within our acceptable range of training costs.
Table 17: The influence of different unzip ratios on the enzyme-substrate affinity regression task and molecule-only tasks.
Unzip Ratio ESAR Molecule-only
[…]
0 0.616 0.554 0.0178 0.0068 74.96 86.46 61.52
1% 0.608 0.558 0.0171 0.0059 77.25 87.13 77.3
5% 0.618 0.552 0.0171 0.0058 78.05 87.65 74